OpenStudy (anonymous):

Choose the Standard Form of the equation y = -one over seven x - 3. Answers one over seven x + y = -3 7x + y = -3 x + 7y = -21 -7x + y = -21

OpenStudy (anonymous):

(1/7)x + y = -3 x + 7y = -21

OpenStudy (anonymous):

What Is It ?

OpenStudy (anonymous):

its x + 7y = -21